OpenStudy (btaylor):

In a classroom with 5 students, is it possible to form 6 distinct groups such that every two groups share exactly 1 student?

OpenStudy (amistre64):

dunno if im reading the question correctly but does 5 choose n = 6 have a solution?

OpenStudy (amistre64):

what does it mean by 6 distinct groups?

OpenStudy (amistre64):

5 choose 3 forms 10 groups abcde abc abd abe acd ace ade bcd bce bde cde also 5 choose 4 gets us 10 groups so i spose it possible to get 6 groups, just need to determine if they fit the rest of the requirements

OpenStudy (btaylor):

Given hint: Suppose you could. Let \(v_1,v_2,\dots,v_6\) be vectors in \(\mathbb{R}^5\) such that \(v_i\) is a vector whose entries consist of 0's and 1's as to encode the \(i\)-th club membership: the \(j\)-th component of \(v_i\) is 1 if student \(j\) is in the \(i\)-th club, and 0 otherwise. What does then the dot product \(v_i \cdot v_j\), for \(i \neq j\), represent? Use that to sow that \(v_1,v_2,\dots,v_6\) must be linear independent vectors in \(\mathbb{R}^5\) (which cannot possibly happen!) as follows: if \(c_1v_1+c_2v_2+\dots + c_6v_6=0\), try expanding \[\left(c_1v_1+c_2v_2+\dots + c_6v_6\right) \cdot \left(c_1v_1+c_2v_2+\dots + c_6v_6\right) = 0\] and conclude that \(c_1=c_2=\dots=c_6=0\). Make sure to use the fact that \(c_1^2+c_2^2+ \dots + c_6^2 + 2\left( c_1c_2+c_1c_3+c_1c_4+\dots+c_5c_6\right)=\left(c_1+c_2+\dots+c_6\right)^2 \) and that \(\left| \left| v_i \right| \right|^2 \ge 1\).

OpenStudy (btaylor):

@amistre64 I think the question is asking if, out of a set of 5 students, you can form 6 groups (of any size, though I'm pretty sure it has to be at least 2) where every set of 2 groups only shares one person.

OpenStudy (btaylor):

@mathmale @ganeshie8 I think you'll enjoy this problem

OpenStudy (amistre64):

so you spose we can have like: ab and bcd as elements to choose from, as opposed to just groups of 2 or 3 or 4 etc

OpenStudy (btaylor):

Yes. There are no given limits on the size of the groups.

OpenStudy (amistre64):

beyond the scope of my ken at the moment ... reality is setting in :)

OpenStudy (anonymous):

yes

OpenStudy (perl):

nice problem

OpenStudy (gorv):

well my view is it is not possible with the only 5 student sharing exactly one sharing